### v2

@@ III.1 $D\bar{B}$ systems
$A$ $A$ $\omega_{1}$ $\omega_{2}$
$A_{a1.1}$ 00 00 $-\delta_{2}$ $-\delta_{2}$
$A_{a1.2}$ 00 00 $-\delta_{1}$ $-\delta_{1}$
$A_{a1.3}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(-3D_{b}+9E_{b})$ $-\delta_{1}$ $-\delta_{2}$
$A_{a1.4}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(-3D_{b}+9E_{b})$ $-\delta_{1}$ $-\delta_{2}$
$A_{a1.5}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ $-\delta_{1}$ 00
$A_{a1.6}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ $-\delta_{2}$ 00
Table 2: The coefficients appearing in the TPE amplitudes (Eqs. (25)-(29)). Note that we have $A_{51}=A_{15}$.
@@
      $A_{c1.1}$       $\frac{1}{16}$       $\frac{1}{16}$       $-\frac{1}{16}$       $-\frac{3}{16}$       $-\frac{3}{16}$       $\frac{3}{16}$       00       00
      $A_{c1.2}$       $\frac{i}{8}$       $-\frac{i}{8}$       00       $-\frac{3i}{8}$       $\frac{3i}{8}$       00       $-\delta_{2}$       00
      $A_{c1.3}$       $\frac{i}{8}$       $-\frac{i}{8}$       00       $-\frac{3i}{8}$       $\frac{3i}{8}$       00       $-\delta_{1}$       00
      $A_{c1.4}$       $\frac{1}{16}$       00       00       $\frac{9}{16}$       00       00       $-\delta_{1}$       $-\delta_{2}$
      $A_{c1.5}$       $\frac{5}{16}$       00       00       $-\frac{3}{16}$       00       00       $-\delta_{1}$       $-\delta_{2}$
@@ III.2 $D\bar{B}^{*}$ systems
×$\times(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (39)
@@
$\mathcal{M}_{c2.2}^{(2)}=$ $~{}i4\frac{g_{2}^{2}}{f^{4}}\{[A_{1}(q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})-A_{5}(q_{0}J_{11}^{S}+2q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})$
$+[A_{1}(q_{0}J_{31}^{S}+J_{34}^{S})-A_{5}(q_{0}J_{21}^{S}+q_{0}J_{31}^{S}+J_{34}^{S})](\epsilon_{2}\cdot\epsilon^{*}_{4})\},$ (42)
$\mathcal{M}_{c2.3}^{(2)}=$ $-i4\frac{g_{2}^{2}}{f^{4}}(d-3)[A_{1}(q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})-A_{5}(q_{0}J_{11}^{S}+2q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})$
$-i4\frac{g_{2}^{2}}{f^{4}}(d-3)\{A_{1}[q_{0}\vec{q}^{2}J_{22}^{S}+\vec{q}^{2}J_{24}^{S}+(2-d)q_{0}J_{31}^{S}+q_{0}\vec{q}^{2}J_{32}^{S}+\vec{q}^{2}J_{33}^{S}+(2-d)J_{34}^{S}]-A_{5}[q_{0}\vec{q}^{2}J_{11}^{S}$
$+(2-d)q_{0}(J_{21}^{S}+J_{31}^{S})+2q_{0}\vec{q}^{2}J_{22}^{S}+\vec{q}^{2}J_{24}^{S}+q_{0}\vec{q}^{2}J_{32}^{S}+\vec{q}^{2}J_{33}^{S}+(2-d)J_{34}^{S}]\}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (43)
$\mathcal{M}_{c2.4}^{(2)}=$ $-i4\frac{g_{1}^{2}}{f^{4}}\{A_{1}[q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}+(1-d)q_{0}J_{31}^{T}+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}+(1-d)J_{34}^{T}]-A_{5}[q_{0}\vec{q}^{2}J_{11}^{T}$
$+(1-d)q_{0}(J_{21}^{T}+J_{31}^{T})+2q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}+(1-d)J_{34}^{T}]\}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (44)
$\mathcal{M}_{c2.5}^{(2)}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A_{1}\{[J_{21}^{B}-\vec{q}^{2}J_{22}^{B}+(d+3)(J_{31}^{B}+J_{42}^{B})-2\vec{q}^{2}J_{32}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})-[\vec{q}^{2}(J_{31}^{B}$
@@
$\mathcal{M}_{c2.6}^{(2)}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{[J_{21}^{B}-\vec{q}^{2}J_{22}^{B}+(d+3)J_{31}^{B}-2\vec{q}^{2}J_{32}^{B}+(d+3)J_{42}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})$
$+[-\vec{q}^{2}J_{21}^{B}+\vec{q}^{4}J_{22}^{B}-(2d+1)\vec{q}^{2}(J_{31}^{B}+J_{42}^{B})+2\vec{q}^{4}J_{32}^{B}+(d+1)(d-2)J_{41}^{B}+\vec{q}^{4}J_{43}^{B}](\epsilon_{2}\cdot\epsilon^{*}_{4})\},$ (46)
$\mathcal{M}_{c2.7}^{(2)}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A_{1}\{[J_{21}^{R}-\vec{q}^{2}J_{22}^{R}+(d+3)(J_{31}^{R}+J_{42}^{R})-2\vec{q}^{2}J_{32}^{R}-\vec{q}^{2}J_{43}^{R}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})-[\vec{q}^{2}J_{31}^{R}$
$-(1+d)J_{41}^{R}+\vec{q}^{2}J_{42}^{R}](\epsilon_{2}\cdot\epsilon^{*}_{4})\},$ (47)
$\mathcal{M}_{c2.8}^{(2)}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{[J_{21}^{R}-\vec{q}^{2}J_{22}^{R}+(d+3)J_{31}^{R}-2\vec{q}^{2}J_{32}^{B}+3J_{42}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})+[-\vec{q}^{2}J_{21}^{R}$
$+\vec{q}^{4}J_{22}^{R}-(2d+1)\vec{q}^{2}(J_{31}^{R}+J_{42}^{R})+2\vec{q}^{4}J_{32}^{R}+(d+1)(d-2)J_{41}^{R}+\vec{q}^{4}J_{43}^{R}](\epsilon_{2}\cdot\epsilon^{*}_{4})\},$ (48)
@@
. $I=1$ $I=0$ $A$ $A$ $\omega_{1}$ $\omega_{2}$ $A_{a2.1}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{3}{4}(D_{a}+E_{a})$ $\delta_{2}$ $\delta_{2}$ $A_{a2.2}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{3}{4}(D_{a}+E_{a})$ 00 00 $A_{a2.3}$ 00 00 $-\delta_{1}$ $-\delta_{1}$ $A_{a2.4}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 00 $-\delta_{1}$ $A_{a2.5}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $-\delta_{1}$ 00 $A_{a2.6}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\delta_{2}$ $-\delta_{1}$ $A_{a2.7}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $-\delta_{1}$ $\delta_{2}$ $A_{a2.(8+9)}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ 00 $\delta_{2}$ $A_{a2.10}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ $-\delta_{1}$ 00
Figure 2: LO contact, NLO contact and NLO TPE diagrams of the process $D\bar{B}^{*}\to D\bar{B}^{*}$. The thin solid, double thin solid, solid, thick solid, and dashed lines stand for the $\bar{B}$, $\bar{B}^{*}$, $D$, $D^{*}$, and a pion, respectively.
Table 4: The coefficients in the TPE amplitudes (Eqs. (41)-(48)). Note that we have $A_{51}=A_{15}$.
@@
       $A_{c2.1}$        $\frac{1}{16}$        $\frac{1}{16}$        $-\frac{1}{16}$        $-\frac{3}{16}$        $-\frac{3}{16}$        $\frac{3}{16}$        00        00
       $A_{c2.2}$        $\frac{i}{8}$        $\frac{-i}{8}$        00        $-\frac{3i}{8}$        $\frac{3i}{8}$        00        00        00
       $A_{c2.3}$        $\frac{i}{8}$        $-\frac{i}{8}$        00        $-\frac{3i}{8}$        $\frac{3i}{8}$        00        $\delta_{2}$        00
       $A_{c2.4}$        $\frac{i}{8}$        $-\frac{i}{8}$        00        $-\frac{3i}{8}$        $\frac{3i}{8}$        00        $-\delta_{1}$        00
       $A_{c2.5}$        $\frac{1}{16}$        00        00        $\frac{9}{16}$        00        00        $-\delta_{1}$        00
       $A_{c2.6}$        $\frac{1}{16}$        00        00        $\frac{9}{16}$        00        00        $-\delta_{1}$        $\delta_{2}$
       $A_{c2.7}$        $\frac{5}{16}$        00        00        $-\frac{3}{16}$        00        00        $-\delta_{1}$        00
       $A_{c2.8}$        $\frac{5}{16}$        00        00        $\frac{-3}{16}$        00        00        $-\delta_{1}$        $\delta_{2}$
@@ III.3 $D^{*}\bar{B}^{*}$ systems
Figure 3: LO contact, LO OPE, NLO contact, NLO OPE, and NLO TPE diagrams of the process $D^{*}\bar{B}^{*}$ to $D^{*}\bar{B}^{*}$. The thin solid, double thin solid, solid, thick solid, and dashed lines stand for the $\bar{B}$, $\bar{B}^{*}$, $D$, $D^{*}$, and a pion, respectively.
@@
At LO, utilizing the Lagrangian presented in Eq. (1) and Eq. (3), the contact and OPE contributions of the scattering process $D^{*}(p_{1})\bar{B}^{*}(p_{2})\to D^{*}(p_{3})\bar{B}^{*}(p_{4})$ depicted in Fig. 3 read
@@
Similar to the definitions (54), we also define
@@
Next, we consider the NLO one-pion-exchange interactions in the $D^{*}\bar{B}^{*}$ systems. The corresponding diagrams are illustrated in Fig. 3, and the amplitudes:
@@
×$\times\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (83)
@@
×$\times\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (84)
@@
×$\times\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (85)
@@
$\mathcal{M}_{c{3.1}}^{(2)}=$ $-4\frac{1}{f^{4}}[A_{1}(q^{2}_{0}J^{F}_{21}+J^{F}_{22})-(A_{15}+A_{51}+A_{5})(q^{2}_{0}J^{F}_{11}+q^{2}_{0}J^{F}_{21}+J^{F}_{22})+A_{5}(q^{2}_{0}J^{F}_{0}+q^{2}_{0}J^{F}_{11})]\mathcal{O}_{1},$ (87)
$\mathcal{M}_{c{3.2}}^{(2)}=$ $-i4\frac{g_{2}^{2}}{f^{4}}\{[A_{1}(q_{0}J^{S}_{22}+J^{S}_{24}+q_{0}J^{S}_{32}+J^{S}_{33})-A_{5}(q_{0}J^{S}_{11}+2q_{0}J^{S}_{22}+J^{S}_{24}+q_{0}J^{S}_{32}+J^{S}_{33})]\mathcal{O}^{a}_{1}+[A_{1}(q_{0}J^{S}_{31}+J^{S}_{34})$
@@
$\mathcal{M}^{(2)}_{c3.3}=$ $~{}i4\frac{g^{2}_{2}}{f^{4}}(d-3)[A_{1}(q_{0}J^{S}_{22}+J^{S}_{24}+q_{0}J^{S}_{32}+J^{S}_{33})-{A}_{5}(q_{0}J^{S}_{11}+2q_{0}J^{S}_{22}+J^{S}_{24}+q_{0}J^{S}_{32}+J^{F}_{33})]\mathcal{O}^{a}_{1}$
$+i4\frac{g^{2}_{2}}{f^{4}}(d-3)\{A_{1}[({q}_{0}\vec{q}^{2}J^{S}_{22}+\vec{q}^{2}J^{S}_{24}+(2-d)q_{0}J^{S}_{31}+q_{0}\vec{q}^{2}J^{S}_{32}+\vec{q}^{2}J^{S}_{33}+(2-d)q_{0}J^{S}_{34})]$
$-{A}_{5}[q_{0}\vec{q}^{2}J^{S}_{11}+(2-d)q_{0}J^{S}_{21}+\vec{q}^{2}J^{S}_{24}+(2-d)q_{0}J^{S}_{31}+q_{0}\vec{q}^{2}J^{S}_{32}+\vec{q}^{2}J^{S}_{33}+(2-d)q_{0}J^{S}_{34}]\}\mathcal{O}_{1},$ (89)
$\mathcal{M}_{c3.4}^{(2)}=$ $~{}i4\frac{g_{1}^{2}}{f^{4}}\{[A_{1}(q_{0}J^{T}_{22}+J^{T}_{24}+q_{0}J^{T}_{32}+J^{T}_{33})-A_{5}(q_{0}J^{T}_{11}+2q_{0}J^{T}_{22}+J^{T}_{24}+q_{0}J^{T}_{32}+J^{T}_{33})]\mathcal{O}^{b}_{1}+[A_{1}(q_{0}J^{T}_{31}+J^{T}_{34})$

### v3

@@ III.1 $D\bar{B}$ systems
AA AA $\omega_{1}$ $\omega_{2}$
$A_{a1.1}$ 0 0 $\delta_{2}$ $\delta_{2}$
$A_{a1.2}$ 0 0 $\delta_{1}$ $\delta_{1}$
$A_{a1.3}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(-3D_{b}+9E_{b})$ $\delta_{1}$ $\delta_{2}$
$A_{a1.4}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(-3D_{b}+9E_{b})$ $\delta_{1}$ $\delta_{2}$
$A_{a1.5}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ $\delta_{1}$ 0
$A_{a1.6}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ $\delta_{2}$ 0
Table 2: The coefficients appearing in the TPE amplitudes [Eqs. (25)-(29)]. Note that we have $A_{51}=A_{15}$.
@@
  $A_{c1.1}$   $\frac{1}{16}$   $\frac{1}{16}$   $\frac{1}{16}$   $\frac{3}{16}$   $\frac{3}{16}$   $\frac{3}{16}$   0   0
  $A_{c1.2}$   $\frac{i}{8}$   $\frac{i}{8}$   0   $\frac{3i}{8}$   $\frac{3i}{8}$   0   $\delta_{2}$   0
  $A_{c1.3}$   $\frac{i}{8}$   $\frac{i}{8}$   0   $\frac{3i}{8}$   $\frac{3i}{8}$   0   $\delta_{1}$   0
  $A_{c1.4}$   $\frac{1}{16}$   0   0   $\frac{9}{16}$   0   0   $\delta_{1}$   $\delta_{2}$
  $A_{c1.5}$   $\frac{5}{16}$   0   0   $\frac{3}{16}$   0   0   $\delta_{1}$   $\delta_{2}$
@@ III.2 $D\bar{B}^{*}$ systems
$\times(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (39)
@@
$\mathcal{M}_{c2.2}^{(2)}=$ $~i4\frac{g_{2}^{2}}{f^{4}}\{[A_{1}(q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})$
$-A_{5}(q_{0}J_{11}^{S}+2q_{0}J_{22}^{S}+J_{24}^{S}$
$+q_{0}J_{32}^{S}+J_{33}^{S})](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})$
$+[A_{1}(q_{0}J_{31}^{S}+J_{34}^{S})$
$-A_{5}(q_{0}J_{21}^{S}+q_{0}J_{31}^{S}+J_{34}^{S})](\epsilon_{2}\cdot\epsilon^{*}_{4})\},$ (42)
$\mathcal{M}_{c2.3}^{(2)}=$ $-i4\frac{g_{2}^{2}}{f^{4}}(d-3)[A_{1}(q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})$
$-A_{5}(q_{0}J_{11}^{S}+2q_{0}J_{22}^{S}+J_{24}^{S}$
$+q_{0}J_{32}^{S}+J_{33}^{S})](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})$
$-i4\frac{g_{2}^{2}}{f^{4}}(d-3)\{A_{1}[q_{0}\vec{q}^{2}J_{22}^{S}+\vec{q}^{2}J_{24}^{S}$
$+(2-d)q_{0}J_{31}^{S}+q_{0}\vec{q}^{2}J_{32}^{S}+\vec{q}^{2}J_{33}^{S}$
$+(2-d)J_{34}^{S}]-A_{5}[q_{0}\vec{q}^{2}J_{11}^{S}$
$+(2-d)q_{0}(J_{21}^{S}+J_{31}^{S})+2q_{0}\vec{q}^{2}J_{22}^{S}+\vec{q}^{2}J_{24}^{S}$
$+q_{0}\vec{q}^{2}J_{32}^{S}+\vec{q}^{2}J_{33}^{S}+(2-d)J_{34}^{S}]\}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (43)
$\mathcal{M}_{c2.4}^{(2)}=$ $-i4\frac{g_{1}^{2}}{f^{4}}\{A_{1}[q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}+(1-d)q_{0}J_{31}^{T}$
$+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}+(1-d)J_{34}^{T}]-A_{5}[q_{0}\vec{q}^{2}J_{11}^{T}$
$+(1-d)q_{0}(J_{21}^{T}+J_{31}^{T})+2q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}$
$+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}+(1-d)J_{34}^{T}]\}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (44)
$\mathcal{M}_{c2.5}^{(2)}=$ $~4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A_{1}\{[J_{21}^{B}-\vec{q}^{2}J_{22}^{B}+(d+3)(J_{31}^{B}+J_{42}^{B})$
$-2\vec{q}^{2}J_{32}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})-[\vec{q}^{2}(J_{31}^{B}$
@@
$\mathcal{M}_{c2.6}^{(2)}=$ $~4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{[J_{21}^{B}-\vec{q}^{2}J_{22}^{B}+(d+3)J_{31}^{B}$
$-2\vec{q}^{2}J_{32}^{B}+(d+3)J_{42}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})$
$+[-\vec{q}^{2}J_{21}^{B}+\vec{q}^{4}J_{22}^{B}-(2d+1)\vec{q}^{2}(J_{31}^{B}+J_{42}^{B})$
$+2\vec{q}^{4}J_{32}^{B}+(d+1)(d-2)J_{41}^{B}+\vec{q}^{4}J_{43}^{B}](\epsilon_{2}\cdot\epsilon^{*}_{4})\},$ (46)
$\mathcal{M}_{c2.7}^{(2)}=$ $~4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A_{1}\{[J_{21}^{R}-\vec{q}^{2}J_{22}^{R}+(d+3)(J_{31}^{R}+J_{42}^{R})$
$-2\vec{q}^{2}J_{32}^{R}-\vec{q}^{2}J_{43}^{R}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})$
$-[\vec{q}^{2}J_{31}^{R}-(1+d)J_{41}^{R}+\vec{q}^{2}J_{42}^{R}](\epsilon_{2}\cdot\epsilon^{*}_{4})\},$ (47)
$\mathcal{M}_{c2.8}^{(2)}=$ $~4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{[J_{21}^{R}-\vec{q}^{2}J_{22}^{R}+(d+3)J_{31}^{R}$
$-2\vec{q}^{2}J_{32}^{B}+3J_{42}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})$
$+[-\vec{q}^{2}J_{21}^{R}+\vec{q}^{4}J_{22}^{R}-(2d+1)\vec{q}^{2}(J_{31}^{R}+J_{42}^{R})$
$+2\vec{q}^{4}J_{32}^{R}+(d+1)(d-2)J_{41}^{R}+\vec{q}^{4}J_{43}^{R}](\epsilon_{2}\cdot\epsilon^{*}_{4})\},$ (48)
@@
. $I=1$ $I=0$ AA AA $\omega_{1}$ $\omega_{2}$ $A_{a2.1}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{3}{4}(D_{a}+E_{a})$ $\delta_{2}$ $\delta_{2}$ $A_{a2.2}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{3}{4}(D_{a}+E_{a})$ 0 0 $A_{a2.3}$ 0 0 $\delta_{1}$ $\delta_{1}$ $A_{a2.4}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 0 $\delta_{1}$ $A_{a2.5}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\delta_{1}$ 0 $A_{a2.6}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\delta_{2}$ $\delta_{1}$ $A_{a2.7}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\delta_{1}$ $\delta_{2}$ $A_{a2.(8+9)}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ 0 $\delta_{2}$ $A_{a2.10}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ $\delta_{1}$ 0
Figure 2: LO contact ($a2$), NLO contact (a2.1a2.10a2.1-a2.10) and NLO TPE (c2.1c2.8c2.1-c2.8) diagrams of the process $D\bar{B}^{*}\to D\bar{B}^{*}$. The thin solid, double-thin solid, solid, thick solid, and dashed lines stand for the $\bar{B}$, $\bar{B}^{*}$, $D$, $D^{*}$, and a pion, respectively.
Table 4: The coefficients in the TPE amplitudes [Eqs. (41)-(48)]. Note that we have $A_{51}=A_{15}$.
@@
  $A_{c2.1}$   $\frac{1}{16}$   $\frac{1}{16}$   $\frac{1}{16}$   $\frac{3}{16}$   $\frac{3}{16}$   $\frac{3}{16}$   0   0
  $A_{c2.2}$   $\frac{i}{8}$   $\frac{-i}{8}$   0   $\frac{3i}{8}$   $\frac{3i}{8}$   0   0   0
  $A_{c2.3}$   $\frac{i}{8}$   $\frac{i}{8}$   0   $\frac{3i}{8}$   $\frac{3i}{8}$   0   $\delta_{2}$   0
  $A_{c2.4}$   $\frac{i}{8}$   $\frac{i}{8}$   0   $\frac{3i}{8}$   $\frac{3i}{8}$   0   $\delta_{1}$   0
  $A_{c2.5}$   $\frac{1}{16}$   0   0   $\frac{9}{16}$   0   0   $\delta_{1}$   0
  $A_{c2.6}$   $\frac{1}{16}$   0   0   $\frac{9}{16}$   0   0   $\delta_{1}$   $\delta_{2}$
  $A_{c2.7}$   $\frac{5}{16}$   0   0   $\frac{3}{16}$   0   0   $\delta_{1}$   0
  $A_{c2.8}$   $\frac{5}{16}$   0   0   $\frac{-3}{16}$   0   0   $\delta_{1}$   $\delta_{2}$
@@ III.3 $D^{*}\bar{B}^{*}$ systems
Figure 3: LO contact ($a3$), LO OPE ($b3$), NLO contact (a3.1a3.2a3.1-a3.20), NLO OPE (b3.1b3.17b3.1-b3.17), and NLO TPE (c3.1c3.13c3.1-c3.13) diagrams of the process $D^{*}\bar{B}^{*}$ to $D^{*}\bar{B}^{*}$. The thin solid, double-thin solid, solid, thick solid, and dashed lines stand for the $\bar{B}$, $\bar{B}^{*}$, $D$, $D^{*}$, and a pion, respectively.
@@
At LO, utilizing the Lagrangian presented in Eqs. (1) and (3), the contact and OPE contributions of the scattering process $D^{*}(p_{1})\bar{B}^{*}(p_{2})\to D^{*}(p_{3})\bar{B}^{*}(p_{4})$ depicted in Fig. 3 read
@@
Similar to the definitions in (54), we also define
@@
Next, we consider the NLO one-pion-exchange interactions in the $D^{*}\bar{B}^{*}$ systems. The corresponding diagrams are illustrated in Fig. 3, and the amplitudes,
@@
$\times\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (83)
@@
$\times\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (84)
@@
$\times\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (85)
@@
$\mathcal{M}_{c{3.1}}^{(2)}=$ $-4\frac{1}{f^{4}}[A_{1}(q^{2}_{0}J^{F}_{21}+J^{F}_{22})$
$-(A_{15}+A_{51}+A_{5})(q^{2}_{0}J^{F}_{11}+q^{2}_{0}J^{F}_{21}+J^{F}_{22})$
$+A_{5}(q^{2}_{0}J^{F}_{0}+q^{2}_{0}J^{F}_{11})]\mathcal{O}_{1},$ (87)
$\mathcal{M}_{c{3.2}}^{(2)}=$ $-i4\frac{g_{2}^{2}}{f^{4}}\{[A_{1}(q_{0}J^{S}_{22}+J^{S}_{24}+q_{0}J^{S}_{32}+J^{S}_{33})$
$-A_{5}(q_{0}J^{S}_{11}+2q_{0}J^{S}_{22}+J^{S}_{24}+q_{0}J^{S}_{32}+J^{S}_{33})]\mathcal{O}^{a}_{1}$
$+[A_{1}(q_{0}J^{S}_{31}+J^{S}_{34})$
@@
$\mathcal{M}^{(2)}_{c3.3}$= $~i4\frac{g^{2}_{2}}{f^{4}}(d-3)[A_{1}(q_{0}J^{S}_{22}+J^{S}_{24}+q_{0}J^{S}_{32}+J^{S}_{33})$
$-{A}_{5}(q_{0}J^{S}_{11}+2q_{0}J^{S}_{22}+J^{S}_{24}+q_{0}J^{S}_{32}+J^{F}_{33})]\mathcal{O}^{a}_{1}$
$+i4\frac{g^{2}_{2}}{f^{4}}(d-3)\{A_{1}[({q}_{0}\vec{q}^{2}J^{S}_{22}+\vec{q}^{2}J^{S}_{24}+(2-d)q_{0}J^{S}_{31}$
$+q_{0}\vec{q}^{2}J^{S}_{32}+\vec{q}^{2}J^{S}_{33}+(2-d)q_{0}J^{S}_{34})]$
$-{A}_{5}[q_{0}\vec{q}^{2}J^{S}_{11}+(2-d)q_{0}J^{S}_{21}+\vec{q}^{2}J^{S}_{24}+(2-d)q_{0}J^{S}_{31}$
$+q_{0}\vec{q}^{2}J^{S}_{32}+\vec{q}^{2}J^{S}_{33}+(2-d)q_{0}J^{S}_{34}]\}\mathcal{O}_{1},$ (89)
$\mathcal{M}_{c3.4}^{(2)}=$ $~i4\frac{g_{1}^{2}}{f^{4}}\{[A_{1}(q_{0}J^{T}_{22}+J^{T}_{24}+q_{0}J^{T}_{32}+J^{T}_{33})$
$-A_{5}(q_{0}J^{T}_{11}+2q_{0}J^{T}_{22}+J^{T}_{24}+q_{0}J^{T}_{32}+J^{T}_{33})]\mathcal{O}^{b}_{1}$
$+[A_{1}(q_{0}J^{T}_{31}+J^{T}_{34})$